# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v1

@@ B.3 Downstream datasets
B.4 Dataset Licenses
The licenses for all datasets are included in the footnotes: fMoW111fMoW license: https://github.com/fMoW/dataset/raw/master/LICENSE, Sentinel-$2$Sentinel-2 license: https://scihub.copernicus.eu/twiki/pub/SciHubWebPortal/TermsConditions/Sentinel_Data_Terms_and_Conditions.pdf, EuroSAT [^3] EuroSAT license: https://creativecommons.org/licenses/by/4.0/, SpaceNet [^4] SpaceNet v1 license: http://creativecommons.org/licenses/by-sa/4.0/, Camelyon17555Camelyon17 license:https://creativecommons.org/publicdomain/zero/1.0/, iWildCam666iWildCam license:https://cdla.dev/permissive-1-0/
@@ Appendix C Environmental Impact
Following [22], we compare the carbon footprint of pre-training using ExPLoRA with domain-specific solutions such as SatMAE. We use the carbon footprint calculator proposed by Lacoste et al. [58]. Our results are in table 12.

### v2

@@ B.5 Downstream datasets
B.6 Dataset Licenses
The licenses for all datasets are included in the footnotes: fMoW111fMoW license: https://github.com/fMoW/dataset/raw/master/LICENSE, Sentinel-$2$Sentinel-2 license: https://scihub.copernicus.eu/twiki/pub/SciHubWebPortal/TermsConditions/Sentinel_Data_Terms_and_Conditions.pdf, EuroSAT [^3] EuroSAT license: https://creativecommons.org/licenses/by/4.0/, SpaceNet [^4] SpaceNet v1 license: http://creativecommons.org/licenses/by-sa/4.0/, Camelyon17555Camelyon17 license:https://creativecommons.org/publicdomain/zero/1.0/, iWildCam666iWildCam license:https://cdla.dev/permissive-1-0/, GlobalWheat [^7] GlobalWheat license:https://opensource.org/licenses/MIT.
@@ Appendix C Environmental Impact
Following Cong et al. (2022), we compare the carbon footprint of pre-training using ExPLoRA with domain-specific solutions such as SatMAE. We use the carbon footprint calculator proposed by Lacoste et al. (2019). Our results are in table 14.